radomes.openskiesproject.org
miajlim
miajlim miajlim
  • 26-05-2024
  • Mathematics
Answered

given that P(A) = 0.82 find P(A')



please give workings out thank you

Verified Answer :

olumideolawoyin
olumideolawoyin olumideolawoyin
  • 26-05-2024

Answer:

0.18

Step-by-step explanation:

Let A = event occuring and A' = event not occuring.

The sum of probabilities of an event occuring and not occuring is 1.

P(A) + P(A') = 1

P(A') = 1 - P(A) since P(A) = 0.82

P(A') = 1 - 0.82

P(A') = 0.18

Answer Link

Related Questions

53x 5 Please Help
Which profitable crop was introduced by john Rolfe
what is a locust who is the first democratic president of rsa
write word equation for the following reactionsHydrochloric acid sodium hydroxideSulphuric acid copper hydroxideNitric acid sodium hydroxideHydrochloric acid ca
What two motions combine to produce orbit
how did emperor franz joseph attempt to restore peace in the austrian empire
The level of competition is an important factor for many people when selecting a physical activity True or false
What is environment
When you add 18 to 14 of a number you get the number itself
Imagine your local police arrest a person suspected of robbing many homes and stores This is an example of how government